OpenStudy (anonymous):

find the rule of : 5,15,45,......,3645

OpenStudy (kira_yamato):

f(x) = 3x

OpenStudy (anonymous):

no thats not it

OpenStudy (zarkon):

\[a_n=5\cdot 3^{n-1}\]

OpenStudy (gw2011):

y=(5)(3)^n-1 where n = the number position in the series. For example, to get the third number (45) in the series, you would do: (5)(3)^3-1 = (5)(3)^2 = (5)(9) = 45